Solve the initial value problem

y"-5y'+6y=2e^x, y(0)=1, y'(0)=1.


1
Expert's answer
2021-11-08T20:25:56-0500

Related homogeneous (complementary) equation


"y''-5y'+6y=0"

Characteristic (auxiliary) equation


"r^2-5r+6=0"

"(r-2)(r-3)=0"

"r_1=2, r_2=3"

The general solution of the homogeneous equation is


"y_h=c_1e^{2x}+c_2e^{3x}"

Find the particular solution of the nonhomogeneous differential equation


"y_p=Ae^x"

Then


"y_p'=Ae^x"

"y_p''=Ae^x"

Substitute


"Ae^x-5Ae^x+6Ae^x=2e^x"

"A=1"

"y_p=e^x"

The general solution of the nonhomogeneous equation is


"y=y_h+y_p"


"y=c_1e^{2x}+c_2e^{3x}+e^x"


Initial conditions

"y(0)=1:"


"1=c_1e^{2(0)}+c_2e^{3(0)}+e^0=>c_1+c_2=0"

"y'(0)=1"


"y'=2c_1e^{2x}+3c_2e^{3x}+e^x"

"1=2c_1e^{2(0)}+3c_2e^{3(0)}+e^0=>2c_1+3c_2=0"


"c_2=-c_1"

"2c_1-3c_1=0"

"c_1=0, c_2=0"

The solution of the given initial value problem is


"y=e^x"